I need to know how to set three different equations that have the solution of 5. Solve your equations for your particular variable to prove the solution is 5.

Answer by stanbon(75887)

Three equations would normally have three variables.
2^5 + 5 + 5*3 = 155
5 - 2*5 + 0 = -20
-5 + 3*5 - 5 = -5
--------------------------
Now put x, y, and z in place of the 5's:
2x + y + 3z = 155
x - 2y + 0 = -20
-x + 3y - z = -5
--------------------------
The solutions should be x = 5, y=5, z=5
Cheers,
Stan H.